OpenStudy (anonymous):

2pi*(4 in) * (5pi/6)/(2pi) = 10pi/3 inches put to a math equation

OpenStudy (anonymous):

\[2\pi*(4\in)*(\frac{ 5\pi }{ 6 }/2\pi=\frac{ 10\pi }{ 3\in }\]